Como resuelvo este sistema por eliminacion?
[ 2x-3y=-2 ]
[ 5x-2y=-5 ]
el corchete que aparece antes del #2 y del #5 no van dividos como los puse aqui, si no que es uno grande que cubre esos dos numeros. Gracias!

Found 2 solutions by richwmiller, TimothyLamb:
Answer by richwmiller(17219)
Answer by TimothyLamb(4379)

2x - 3y = -2
5x - 2y = -5
---
(5)2x - (5)3y = -(5)2
(-2)5x - (-2)2y = -(-2)5
---
add (sumar)
10x - 15y = -10
-10x + 4y = 10
---
-11y = 0
y = 0
---
answer (respuesta):
y = 0
5x - 2(0) = -5
5x = -5
x = -1
---
(-1, 0)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php